# GLOV: Guided Large Language Models as Implicit Optimizers for Vision Language Models

### v1

@@ 3.2 Meta-Prompt
In-context examples serve to bootstrap the LLM to the type of output that is expected. In each optimization step, we task the LLM to provide us with $10$ candidate solutions (prompts). Each prompt is evaluated w.r.t. a fitness function to obtain a (classification) score. We keep a global history of the prompts (and associated fitness) generated during all the previous optimization steps and at each optimization step $i$, the newly generated prompts and all the previous prompts are ranked according to their respective fitness score. For the next optimization step $i+1$, $top_{k}$, and $bottom_{k}$ prompts (we choose $k=5$) are selected as in-context demonstrations and plugged into the meta-prompt together with their respective accuracies. The intuition behind keeping a global history of prompts is to provide the LLMs with long-term knowledge about what types of prompts have been effective for the VLM so that it can model its responses according to them.
@@ 3.3 Steering the LLM Generation Process
For each new token produced in the subsequent optimization iteration, the difference between $H_{+}$ and $H_{-}$ is added autoregressively to the embeddings of each generated token [^1] We also experiment with several alternatives for adding the offset vector in the ablations Section 4.3, however, we find that adding the offset to only the last token performs best.. Let $\vec{p}_{n}$ denote the new token appended to the (meta) prompt, then the updated sentence embedding $H_{n}$ is given by
@@
where $\alpha$ is the scaling factor and is chosen via grid search. This process is repeated until the maximum number of tokens is achieved for each prompt. In total we prompt the LLM (at each iteration) to provide us with $10$ prompt templates for CLIP and $5$ for LLaVa to reduce the computation efforts. In an optimization run $p_{+}$ is always the prompt with the best accuracy w.r.t the fitness and $p_{-}$ is set to be the prompt with the second-best accuracy. Since, we compute a form of the gradient-like differential between averages of token hidden states, intuitively trying to identify a characteristic of task-specific improvement. Thus, the intuition behind computing the differential between the best and the second best (in terms of fitness) is to make it between points closest to the maximal value of the objective – which is a common mathematical intuition. Furthermore, $p_{+}$ and $p_{-}$ are only updated when a new prompt with higher accuracy is found. This ensures that the guidance signal does not alter in each iteration, resulting in more stable optimization.
An important design choice in GLOV is the method adopted to calculate the sentence embeddings. Some works, e.g., Jiang et al. (2023) hint that the decoder-based LLMs are not suitable for obtaining the sentence embeddings. We ablate our proposed method of obtaining sentence embeddings (equation 3) in Section 4.3 and find it to provide strong results (while linear probing the embeddings from the middle layers of the LLM) on common natural language classification tasks, hinting that our sentence embeddings can capture semantically meaningful information from the prompts.
4 Experimental Evaluations
@@
CLIP (S-TEMP) (Radford et al., 2021) 61.9 54.8 91.4 65.4 40.3 28.2 64.0 81.3
CLIP (DS-TEMP) (Radford et al., 2021) 63.3 56.0 89.9 67.9 42.1 30.2 66.6 83.2
LLM-OPT (Liu et al., 2024b) 62.8 55.6 92.3 67.5 41.9 28.1 67.0 78.1
@@
CLIP (S-TEMP) (Radford et al., 2021) 60.2 40.2 77.6 18.1 62.1 60.4 54.1 35.8
CLIP (DS-TEMP) (Radford et al., 2021) 59.9 42.4 79.2 19.4 61.7 62.3 57.2 45.8
LLM-OPT (Liu et al., 2024b) 60.2 41.7 79.2 17.7 60.9 60.9 54.4 45.0
@@
In this section, we first provide a brief overview of the datasets we use for evaluating our GLOV, then provide an overview of the different baselines and state-of-the-art methods we compare to, later discuss the implementation details and finally conclude with a discussion on the results.
4.1 Evaluation Settings
Datasets: We extensively evaluate our GLOV on $16$ object recognition datasets belonging to widely different domains. These domains can be narrowed down to datasets containing commonly occurring natural categories: ImageNet (Deng et al., 2009), ImageNetV2 (Recht et al., 2019), Caltech101 (Fei-Fei et al., 2004), fine-grained classification datasets containing different task-specific images: Oxford Flowers (Nilsback & Zisserman, 2008), Standford Cars (Krause et al., 2013), Oxford Pets (Parkhi et al., 2012), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Food-101 (Bossard et al., 2014), FGVC-Aircraft (Maji et al., 2013). Dataset used for scene classification: SUN397 (Xiao et al., 2010), action recognition dataset: UCF101 (Soomro et al., 2012). Datasets consisting of out-of-distribution images: ImageNet-(R)endition (Hendrycks et al., 2021a), ImageNet-(A)dversarial (Hendrycks et al., 2021b), ImageNet-(S)ketch (Wang et al., 2019) and also datasets which contain images taken from a satellite or an aerial view: EuroSAT (Helber et al., 2018) and RESISC45 (Cheng et al., 2017).
Baselines: We compare to the following baselines and state-of-the-art methods:
CLIP (Radford et al., 2021) denotes the zero-shot classification scores obtained by using the simple ‘{a photo of a <class name>}’ template (S-TEMP) and dataset-specific templates (DS-TEMP [^2] https://github.com/openai/CLIP/blob/main/data/prompts.md). LLaVa-OV represent results obtained by using a base prompt [^3] The prompt used to obtain the base results is: Describe the category present in this image briefly and also identify the name of the category present, which was able to match or surpass accuracy reported by Zhang et al. (2024).. LLM-OPT (Liu et al., 2024b) proposes to find suitable text prompts for the downstream datasets by iterative refinement through an LLM. However, their method relies only on the in-context examples without explicit guidance [^4] Note that this method is not suitable for application on the encoder-decoder models because it requires a memory bank of templates, which are not available a-priori for these models.. Meta-prompt (Mirza et al., 2024) propose a method for improving the visual recognition performance of dual-encoder models by generating category-level VLM prompts. Here we extend its evaluations to dual-encoder models [^5] The original publication generates category-level prompts for the dual-encoder models, which is a different setting than the one studied in our work.. For completeness, we also compare with the gradient-based (few-shot) learning method CoOp (Zhou et al., 2022) in the Appendix Table 5.
Implementation Details: To report the results for each dataset we use the test splits provided by Zhou et al. (2022). All the baselines are also implemented in the same framework. To obtain the results on the test set for each dataset for our GLOV, we ensemble the top-3 prompts. These prompts are chosen with regard to the best-performing prompts on the 1-shot train set at a certain iteration during the optimization. For our GLOV we use Llama-3 (Touvron et al., 2023a) from Hugging Face, whereas, for LLM-OPT we keep consistent with their original implementation regarding all details. The best-performing prompts for LLM-OPT are also chosen w.r.t the 1-shot train set. We set the maximum number of optimization iterations to $100$ (with $10$ candidate solutions at each iteration) for the experiments with CLIP and $50$ (with $5$ candidate solutions) for LLaVa-OV, except for datasets containing $1000$ classes (e.g., ImageNet), where we set the maximum number of iterations to $25$ to save computation time. In general, the experiments with CLIP can run on a single NVIDIA 3090 (24GBs) GPU, and the experiments with LLaVa fit on an A40 (48GBs) GPU or similar.
@@ 4.1 Evaluation Settings
Table 2: Results on encoder-decoder VLM. Top-1 accuracy (%) for $16$ datasets obtained by employing the LLaVa (One Vision) (Li et al., 2024). LLaVa (OV) refer to the results obtained by using a generic prompt, while Meta-Prompt refer to the results obtained by obtaining the prompts through Mirza et al. (2024).
@@ 4.2 Results
We evaluate our GLOV extensively on $16$ diverse datasets. In Table 1 we list the results by employing the CLIP ViT-B/32 from OpenAI. We observe that our GLOV achieves better accuracy on all the datasets evaluated. For example, as compared to CLIP, when using the simple prompt template, our vanilla GLOV provides an average gain of $2.5\%$, with up to 11.0% gains on EuroSAT. Similarly, the gains increase even further with our proposed guidance scheme. We observe gains of up to $15\%$ ($3.8\%$ on average). On the large-scale ImageNet dataset, we observe gains of $2.5\%$, which shows that our guidance can help to find prompts that are generalizable across diverse categories of ImageNet. On the other out-of-distribution (OOD) ImageNet variants, our GLOV is also able to show consistent improvements. We also compare with the CLIP classifier constructed by ensembling the hand-crafted templates provided by OpenAI2 and find that the prompts discovered through our GLOV can even improve upon these results. For example, on the large-scale ImageNet dataset, the CLIP classifier built by ensembling the top-3 performing prompts on the train set can provide a gain of $1.2\%$, while on average the performance improvements is $1.5\%$, with up to $5.0\%$ gains on the EuroSAT dataset. It is also important to point out that the prompts provided by CLIP (Radford et al., 2021) are chosen w.r.t the accuracy on the test set (Liu et al., 2024b), whereas, our GLOV searches for prompts by only having access to $1$-shot training data highlighting the generalization ability of our GLOV. Furthermore, as compared to ensembling $80$ CLIP templates, our results are obtained by only ensembling the top-3 prompts, highlighting redundancy in CLIP prompts.
@@
In Table 1, we also compare our GLOV with LLM-OPT (Liu et al., 2024b). Even our vanilla prompting method (without guidance) is on average $1.4\%$ better than LLM-OPT. This highlights that our meta-prompt is better suited to the task of prompt search. Our meta-prompt consists of task-specific and long-term knowledge of the LLM’s responses about what it has generated in all the previous iterations, whereas, LLM-OPT’s prompt does not contain long-term dependencies. Furthermore, we are providing the absolute accuracy associated with each of the in-context examples, which helps the LLM with fine-grained knowledge about the effectiveness of the prompt. On the other hand, the prompt used by LLM-OPT (Liu et al., 2024b), naïvely instructs the LLM to provide better prompts, with only providing good and bad prompts. From the results, we also observe that our proposed guidance methodology further helps to obtain better results by obtaining $2.7\%$ improvement on average, and with $1.7\%$ improvements on the large-scale ImageNet. These results highlight the effectiveness of our GLOV, which is further strengthened by our novel guidance mechanism.
@@
In Table 2 we list the detailed results by employing the LLaVa-OV-7B model. Due to the generative nature of these models, recent works (Geigle et al., 2024, Zhang et al., 2024) have highlighted the difficulty faced in evaluating these models for fine-grained visual recognition. Specifically, Zhang et al. (2024) proposes to fine-tune the models to improve their visual recognition performance. However, in our work, we find that for these models, the fine-grained visual recognition performance can be greatly enhanced by finding the optimal prompt (without requiring gradient-based learning). We observe that the solutions discovered by our GLOV can significantly close the gap with the dual-encoder models. For example, we observe up to $57.5\%$ improvement ($21.5\%$ on average over $16$ datasets) as compared to vanilla LLaVa-OV. We also observe that our proposed guidance scheme has a significant impact on the results. For example, after obtaining the prompt by directing the LLM towards the better solutions discovered, we observe a significant average improvement (over $16$ datasets) over the vanilla GLOV of $2.5\%$ and notably on the large-scale ImageNet dataset GLOV-guidance obtains a healthy improvement of $4.9\%$. These results display the effectiveness of our proposed embedding space guidance schema.
These results show the quantitative benefits of our GLOV. We also visualize the evolution of accuracy on the train set as the optimization process proceeds in Figure 1 for the ImageNet dataset. We observe that as the prompt optimization proceeds, our GLOV shows a consistent increase in accuracy, and also our proposed guidance fares better than the vanilla GLOV. These plots indicate that the LLM gradually starts to understand the type of language preferred by the downstream VLM (while only being interfaced with a fitness function) and our proposed guidance helps to provide it a direction, which is followed by the LLM to continuously discover solutions that improve the downstream vision task. We delegate the actual (best) prompts found, the evolution of prompts, and the optimization evolution for other datasets to Appendix Section B.

### v5

@@ 3.2 LLM Prompt
In-context examples serve to bootstrap the LLM to the type of output that is expected. In each optimization step, we task the LLM to provide us with $10$ candidate solutions (prompts). Each prompt is evaluated w.r.t. a fitness function to obtain a (classification) score. We keep a global history of the prompts (and associated fitness) generated during all the previous optimization steps and at each optimization step $i$, the newly generated prompts and all the previous prompts are ranked according to their respective fitness score. For the next optimization step $i+1$, $top_{k}$, and $bottom_{k}$ prompts (we choose $k=5$) are selected as in-context demonstrations and plugged into the prompt together with their respective accuracies. The intuition behind keeping a global history of prompts is to provide the LLMs with long-term knowledge about the types of prompts that have been effective for the VLM so that it can model its responses according to them.
@@ 3.3 Steering the LLM Generation Process
For each new token produced in the subsequent optimization iteration, the difference between $H_{+}$ and $H_{-}$ is added autoregressively to the embeddings of each generated token [^2] We also experiment with several alternatives for adding the offset vector in the appendix Section B.1, however, we find that adding the offset to only the last token performs best.. Let $\vec{p}_{n}$ denote the new token appended to the LLM prompt, then the updated sentence embedding $H_{n}$ is given by
@@
where $\alpha$ is the scaling factor and is automatically optimized by GLOV on the held-out training set, by an $\alpha$ sweep. This process is repeated until the maximum number of tokens is achieved for each prompt. In total we prompt the LLM (at each iteration) to provide us with $N$ prompt templates. In an optimization run $p_{+}$ is always the prompt with the best accuracy w.r.t the fitness and $p_{-}$ is set to be the prompt with the second-best accuracy. Since, we compute a form of the gradient-like differential between averages of token hidden states, intuitively trying to identify a characteristic of task-specific improvement. Thus, the intuition behind computing the differential between the best and the second best (in terms of fitness) is to obtain it between points closest to the maximal value of the objective – which is a common mathematical intuition. Furthermore, $p_{+}$ and $p_{-}$ are only updated when a new prompt with higher accuracy is found. This ensures that the guidance signal does not alter in each iteration, resulting in more stable optimization. The detailed algorithm is laid out in Algorithm 1.
We ablate our proposed method of obtaining sentence embeddings (equation 3) in Section 4.3 and find it to provide strong results (while linear probing the embeddings from the middle layers of the LLM) on common natural language classification tasks, hinting that our sentence embeddings can capture semantically meaningful information from the prompts. This analysis also motivates our choice of relying on the middle layers for extracting guidance in GLOV.
@@
CLIP (S-TEMP) 61.9 54.8 91.4 65.4 40.3 28.2 64.0 81.3
CLIP (DS-TEMP) 63.3 56.0 89.9 67.9 42.1 30.2 66.6 83.2
LLM-OPT 62.8 55.6 92.3 67.5 41.9 28.1 67.0 78.1
@@
CLIP (S-TEMP) 60.2 40.2 77.6 18.1 62.1 60.4 54.1 35.8
CLIP (DS-TEMP) 59.9 42.4 79.2 19.4 61.7 62.3 57.2 45.8
LLM-OPT 60.2 41.7 79.2 17.7 60.9 60.9 54.4 45.0
@@
Table 2: Results on encoder-decoder VLM. Top-1 accuracy (%) for $16$ datasets obtained by employing the LLaVA (One Vision) (Li et al., 2024). LLaVA-OV refer to the results obtained by using a generic prompt.
4 Experimental Evaluations
In this section, we first list the datasets employed for evaluating our GLOV, then provide an overview of the different baselines and state-of-the-art methods we compare to, later discuss the implementation details and finally conclude with a discussion of the results.
4.1 Evaluation Settings
Image Recognition Datasets: We extensively evaluate our GLOV on $16$ object recognition datasets belonging to widely different domains. These domains can be narrowed down to datasets containing commonly occurring natural categories: ImageNet (Deng et al., 2009), ImageNetV2 (Recht et al., 2019), Caltech101 (Fei-Fei et al., 2004), fine-grained classification datasets containing different task-specific images: Oxford Flowers (Nilsback & Zisserman, 2008), Standford Cars (Krause et al., 2013), Oxford Pets (Parkhi et al., 2012), Describable Textures Dataset (DTD) (Cimpoi et al., 2014), Food-101 (Bossard et al., 2014), FGVC-Aircraft (Maji et al., 2013). Dataset used for scene classification: SUN397 (Xiao et al., 2010), action recognition dataset: UCF101 (Soomro et al., 2012). Datasets consisting of out-of-distribution images: ImageNet-(R)endition (Hendrycks et al., 2021a), ImageNet-(A)dversarial (Hendrycks et al., 2021b), ImageNet-(S)ketch (Wang et al., 2019) and also datasets which contain images taken from a satellite or an aerial view: EuroSAT (Helber et al., 2018) and RESISC45 (Cheng et al., 2017).
Safety Datasets: We evaluate our GLOV on two recently proposed safety benchmarks for VLMs. MM-Safety-Bench (Liu et al., 2024c) contains $3$ splits based on the image source (typography, images generated from diffusion models, and a combination), and the text instructions are divided into $13$ categories (e.g., fraud, hate speech, etc.). VLGuard (Zong et al., 2024) contains $3$ splits: unsafes, safe-unsafes, safe-safes. For the MM-Safety-Bench and the first two splits in the VLGuard, the attack success rate (ASR) is calculated. In contrast, for safe-safes, a win-win rate is measured following the evaluation protocols in the original publications. More details are provided in the Appendix Section C.
Baselines: We compare to the following baselines and state-of-the-art methods: CLIP (Radford et al., 2021) denotes the zero-shot classification scores obtained by using the simple ‘{a photo of a <class name>}’ template (S-TEMP) and dataset-specific templates (DS-TEMP [^3] https://github.com/openai/CLIP/blob/main/data/prompts.md). For image classification, LLaVA-OV represent results obtained by using a base prompt [^4] The prompt utilized to get the base results is: Describe the category present in this image briefly and also identify the name of the category present, which was able to match or surpass accuracy reported by (Zhang et al., 2024).. LLM-OPT (Liu et al., 2024b) proposes to find suitable text prompts for the downstream datasets by iterative refinement through an LLM. However, their method relies only on the in-context examples without explicit guidance [^5] Note that this method is not suitable for application on the encoder-decoder models because it requires a memory bank of templates, which are not available a-priori for these models.. For evaluations on the safety benchmarks, we incorporate another state-of-the-art VLM, Molmo (Deitke et al., 2024). For completeness, we also compare with the gradient-based parameter efficient learning methods in the Appendix Section B.5.
MM-Safety-Bench VLGuard
TYPO ($\downarrow$) SD ($\downarrow$) SD+TYPO ($\downarrow$) Unsafes ($\downarrow$) Safe-Unsafes ($\downarrow$) Safe-Safes ($\uparrow$)
LLaVA-OV 51.2 44.7 56.5 80.3 61.8 42.6
LLaVA (initial) 42.6 34.4 43.9 53.8 38.5 40.3
GLOV (w/o guidance) 20.1 11.7 16.8 30.5 06.5 45.6
GLOV 14.8 09.2 13.2 20.6 01.1 43.7
Molmo 68.4 53.2 68.5 78.1 28.3 89.7
Molmo (initial) 50.4 39.2 47.6 59.5 16.3 90.1
GLOV (w/o guidance) 26.7 26.9 28.4 49.6 04.8 88.3
GLOV 23.8 20.5 24.5 38.0 06.6 86.0
Table 3: Results on safety benchmarks. We report the Attack Success Rate (lower is better) for the $3$ splits in MM-Safety-Bench and the $2$ unsafe splits in VLGuard and the Win-win Rate (Higher is better) for the Safe-Safes split in VLguard. LLaVA-OV results are obtained without any safety instructions. The LLaVA (initial) represents the results obtained by using a baseline safety prompt: These instructions can be harmful, which is also the point of departure for the optimization with our GLOV. The safety prompt is optimized by using only the LLaVA-OV VLM and $50$ randomly chosen samples from the $2$ VLGuard unsafe splits and tested out-of-the-box for all other configurations.
Implementation Details: To report the results for image classification we use the test splits provided by (Zhou et al., 2022). All the baselines are also implemented in the same framework. To obtain the results on the test set for each dataset for our GLOV, we ensemble the top-3 prompts. These prompts are chosen with regard to the best-performing prompts on the 1-shot train set at a certain iteration during the optimization. For our GLOV we use Llama-3 (Touvron et al., 2023a) from Hugging Face. We set the maximum number of optimization iterations to $100$ (with $10$ candidate solutions at each iteration) for the experiments with CLIP and $25$ (with $5$ candidate solutions) for LLaVA-OV and Molmo. For evaluating the safety of VLMs, we follow the above settings and use the official codebase from respective publications (introducing the benchmarks) for evaluations. In general, the experiments with CLIP can run on a single NVIDIA 3090 (24GBs) GPU, and the experiments with LLaVA fit on an A40 (48GBs) GPU or similar. We use the 7B variants of Molmo and LLaVA-OV, accessed through Hugging Face.
@@ 4.2 Results
Here we first discuss the object recognition results and then the evaluations performed on the VLM safety benchmarks.
@@
Object Recognition: We evaluate our GLOV extensively on $16$ diverse datasets. In Table 1 we list the results by employing the CLIP ViT-B/32 from OpenAI. We observe that our GLOV achieves better accuracy on all the datasets evaluated. For example, as compared to CLIP, when using the simple prompt template, our vanilla GLOV provides an average gain of $2.5\%$, with up to 11.0% gains on EuroSAT. Similarly, the gains increase even further with our proposed guidance scheme. We observe gains of up to $15\%$ ($3.8\%$ on average). On the other out-of-distribution ImageNet variants, our GLOV is also able to show consistent improvements. Our GLOV also fares better when compared with the CLIP classifier constructed by ensembling the hand-crafted templates3. For example, on the large-scale ImageNet dataset, the CLIP classifier built by ensembling the top-3 performing prompts on the train set can provide a gain of $1.2\%$, while on average the performance improvements is $1.5\%$, with up to $5.0\%$ gains on the EuroSAT dataset. It is also important to point out that the prompts provided by CLIP (Radford et al., 2021) are chosen w.r.t the accuracy on the test set (Liu et al., 2024b), whereas, our GLOV searches for prompts by only having access to $1$-shot training data highlighting the generalization ability of our GLOV.
@@
In Table 1, we also compare our GLOV with LLM-OPT (Liu et al., 2024b). Even our vanilla prompting method (without guidance) is on average $1.4\%$ better than LLM-OPT, while the proposed guidance scheme provides $2.7\%$ improvement on average, with $1.7\%$ improvements on the large-scale ImageNet. This highlights that our prompting methodology coupled with the guidance scheme is better suited to the task of prompt search. Our prompt consists of task-specific and long-term knowledge of the LLM’s responses about what it has generated in all the previous iterations, whereas, LLM-OPT’s prompt does not contain long-term dependencies. Furthermore, the proposed guidance scheme induces a notion of goodness to the language generation, that is reflected in the downstream results. On the other hand, the prompt used by LLM-OPT (Liu et al., 2024b), naïvely instructs the LLM to provide better prompts, with only providing good and bad prompts, without any explicit guidance, making it challenging for the LLM to bias its outputs towards preferred langugage.
@@
In Table 2 we list the detailed results by employing the LLaVA-OV-7B model. Zhang et al. (2024) find that generative models have extremely low fine-grained visual recognition performance and they improve it via finetuning. In our work, we find that for these models, the fine-grained visual recognition performance can be greatly enhanced by finding the optimal prompt (without requiring gradient-based learning). We observe that the solutions discovered by our GLOV can significantly close the gap with the dual-encoder models. For example, we observe up to $57.5\%$ improvement ($21.5\%$ on average over $16$ datasets) as compared to vanilla LLaVA-OV. We also observe that our proposed guidance scheme has a significant impact on the results. The guidance further provides $2.5\%$ average improvement over the vanilla GLOV, and notably on the large-scale ImageNet dataset the improvement is $4.9\%$.
Classes 1000 50 20 15 10 LLaVA-OV
Accuracy 51.7 49.9 47.0 46.6 44.8 36.5
Classes 200 20 15 10 5 LLaVA-OV
Accuracy 77.6 72.6 71.1 68.6 54.3 52.1
Table 4: Sub-shot setting. Top-1 Accuracy (%) by using LLaVA-OV for ImageNet (left) and ImageNet-R (right) for different number of classes used for GLOV optimization. For ImageNet, using $10$ classes would mean using only 10 labeled samples (from $10$ classes) from the ImageNet dataset instead of $1000$ (from 1000 classes) for our fitness calculation.
DTD EuroSAT IN-R IN-A UCF-101
CLIP 40.2 35.8 65.4 28.2 60.4
GLOV [Llama-3-7b] 42.6 50.8 68.5 32.5 63.8
GLOV [Llama-3.1-70b] 44.5 54.0 69.7 33.3 64.7
Table 5: Scaling LLM size. Top-1 Accuracy (%) for datasets by using different variants of Llama.
Safety Enhancement: In Table 3 we provide results on two recent benchmarks: MM-Safety-Bench (Liu et al., 2024c) and VLGuard (Zong et al., 2024). These benchmarks evaluate the harmful nature of responses from the VLMs given an image and a corresponding instruction. We find that our GLOV can greatly enhance the safety of the present VLMs by finding a suitable safety instruction – prepended to the original text instruction. For example, we observe that our GLOV can reduce the harmful responses by up to $60.7\%$ ($60.2\%$ on average) while compared to the base VLM (LLaVA-OV) - prompted without any safety instruction on the VLGuard dataset. We also observe no loss of win-win rate on the Safe-Safes split, which evaluates the models’ ability to answer safe questions. We also evaluate the recent Molmo VLM and find similar trends. The detailed results for the MM-Safety-Bench are provided in the Appendix Section D. Note that we only optimize for the safety instructions on $50$ samples randomly chosen from the VLGuard unsafe splits (with LLaVA-OV) and evaluate on all other configurations evaluated in Table 3. These results highlight that the optimized safety instructions are general and they transcend models and benchmarks.
Ideally, the safety instructions should not hurt the model performance on general tasks. To evaluate this, we test the VLMs by prepending the safety instructions to the questions in SEED (Fang et al., 2021) and MME (Yin et al., 2023) benchmarks (c.f., ablations Table 6). Interestingly, we find that in some cases the safety instruction enhances the performance of the model (by up to 23.3 points) on these general-purpose benchmarks, possibly by forcing the model to look closely at the instructions. To avoid clutter, we delegate the actual (best) prompts found, the evolution of prompts, and the optimization evolution for datasets to the Appendix Section E.